OpenStudy (anonymous):

What is the slope of a line that passes through the coordinates six comma four and the coordinates negative one comma three?

OpenStudy (raden):

use the formula : m = (y2 - y1)/(x2 - x1) with m is the slope

OpenStudy (anonymous):

I didn't get the answer I got the wrong answer

OpenStudy (raden):

given 2 coordinat points (x1, y1) = (6,4) and (x2, y2) = (-1,3) distritribute that points into formula above

OpenStudy (anonymous):

ok thanks

OpenStudy (raden):

what do you get ?

OpenStudy (anonymous):

I got negative 7 is it right?

OpenStudy (raden):

m = (y2 - y1)/(x2 - x1) m = (3 - 4)/(-1 - 6) = -1/-7 = 1/7

OpenStudy (anonymous):

i did it over and got that lol thanks!

OpenStudy (raden):

welcome :)